OpenStudy (anonymous):

Someone help me please . Write an equation to model the given situation. The area of a rectangle is 48 square centimeters. The length of the rectangle is 8 centimeters longer than the width.

OpenStudy (anonymous):

Let's use l to denote length and w to denote width. What is the formula for computing the area of a rectangle given its length and width?

OpenStudy (anonymous):

\[Area = l \times w\]

OpenStudy (anonymous):

\[l = w + 8\]

OpenStudy (anonymous):

Length is always longer than the width. In this case, the length is exactly 8cm longer than the width.

OpenStudy (anonymous):

We can then express the area of this rectangle in either length only or width only.

OpenStudy (anonymous):

\[Area = l \times (l-8) = w \times (w+8) = 48\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Please feel free to medal if you believe this solution was helpful.